# Slow is Fast! Dissecting Ethereum’s Slow Liquidity Drain Scams

### v2

@@ 9 Conclusion
This paper introduces SLID scams as a stealthy and financially devastating threat to the DeFi ecosystem. By combining heuristic-based detection with machine learning, we identified 3,117 SLID pools, exposing \$103.2M in cumulative losses. Our ML model achieves early detection, enabling timely intervention to protect investors. We have responsibly disclosed our findings to the affected DEXs. We hope this study raises community awareness to prevent such gradual, daily losses. “Slow” does not mean safe; it can silently drain your wallet.
Appendix A Additional Related Work
@@
Liquidity pool, DEX and DeFi. Over the past five years, the rise of DeFi and DEXs has attracted millions of users and billions in inflows, leading to numerous research studies [64, 65, 66, 56, 9, 67, 68, 69]. Hendrik et al. [64] examined the opportunities and challenges for DeFi and its products in this mass adoption era. Fan et al. [66] conducted a comprehensive study on cryptocurrency trading activities and methods for measuring market valuation. Jianlei et al. [56] investigated the DEXs roles and their impact on the development of the DeFi ecosystem. Saleh et al. [67] analyzed the effects of liquidity on Uniswap for liquidity providers and conducted estimations of fees and impermanent loss. Similarly, Lioba et al. [68] developed a theoretical model to analyze the risks and returns faced by liquidity providers under various circumstances.
@@
Cryptocurrency scam detection. Various research efforts have focused on analyzing DEX scams, such as rug pull [24, 37, 70, 71] and honeypot [25, 48, 72, 73]. For instance, Bruno et al. [24] utilized ML algorithms to identify rug pull and scam-likely tokens. Cernera et al. [37] conducted a longitudinal analysis of the lifetime of rug pull liquidity pools on Ethereum and Binance Smart Chain. Rundong et al. [48] introduced various types of honeypot traps on DEXs along with their respective detection methods. Similarly, Christof et al. [25] conducted a comprehensive honeypot analysis with a proposed symbolic-execution-based solution. In addition, the scams originating from traditional finance, such as pump-and-dump schemes [74, 75] and Ponzi schemes [44, 76], have also been studied to provide warnings to investors. Several studies [77, 78, 79, 80] have suggested general fraud detection models that identify anomalous transactions or users through graph neural networks. However, among all the studies in this field, the SLID scam has yet to be discovered.
Appendix B Pool Owner Financial Guarantee in SLID
@@
Let the liquidity pool’s initial deposit (by the owner) be $x$ units of the paired token and $y$ units of the base token. From the formula of the AMM model from Uniswap [34], the relationship between these tokens in the pool is reflected by:
@@
where $k$ is a constant determined by the initial liquidity and remains unchanged. With the mentioned setup, the pool’s owner initially monopolizes the paired token supply, which means no paired tokens $x$ exist outside the pool before the pool’s deployment. Any investor/user that is investing/interacting with the pool must first buy the paired token from the pool by swapping it with base tokens. Let $\Delta x$ represent the total amount of the paired token bought by users or distributed as incentives (for users that deposit a pair of tokens for liquidity providers). The formula will be rewritten to represent the pool’s state any moment after deployment:
@@
where $\Delta y$ is the change of base token amount in the pool due to the change of paired token, and x’/y’ is the amount of paired/base token in the pool after the change compared to the initial deposit amount (from the owner). Under the AMM model, the constant $k$ must not change.
In the worst scenario where no investor interacts with the pool ($\Delta x=0$) or all investors sell back their paired token ($=\Delta x$, since the monopoly of paired tokens from the owner made the limit of paired token that investor possess totally at any time cannot be larger than $\Delta x$) to the pool, the base token possessed by the owner afterward is computed by:
$y^{\prime}=\frac{k}{x^{\prime}}=\frac{x\times y}{x+\Delta x-\Delta x}=y$ (4)
This proves that in the worst-case scenario, where the pool has no investor or all investors that previously bought (and use the bought paired token to deposit for incentives) sell back all their received paired tokens to the pool, the base token balance controlled by the owner will still equal to $y$ which is the initial investment of the owner to the pool as initial liquidity. Hence, with this setup, the pool owner is guaranteed to maintain their original investment of base tokens $y$, which contain real value, under any sequence of investor interactions after the pool’s deployment.
Appendix C Computing Equation
@@
$\begin{split}\textit{Profit\_realized}=&\left(\sum_{i=1}^{m}Sell_{i}+\sum_{j=1}^{n}Withdraw_{j}\right)\\ &-\left(\sum_{k=1}^{o}Buy_{k}+\sum_{l=1}^{p}Deposit_{l}+\text{GAS}\right)\end{split}$ (5)
@@
where the owner’s investment activity (which adds assets in base tokes to the pool) includes $Buy_{k}$ and $Deposit_{l}$, representing the asset value (in USD) added to their liquidity pool during the $k^{\text{th}}$ buy and $l^{\text{th}}$ deposit orders, respectively. In contrast, the owner’s realized return consists of $Sell_{i}$ and $Withdraw_{j}$, which are the asset values in base token (in USD) that the owner removed from the pool during the $i^{\text{th}}$ sell and $j^{\text{th}}$ withdraw orders, respectively. The metric also accounts for the total network gas fee (GAS) used for pool deployment and all DEX activities performed by the owner. Since the equation focused on computing realized profit, all of the variables are only around the owner’s balance change of the base token (which holds the real value) through their activity in the pool (which also involves paired token as well).
@@
Unrealized profit computing equation.  To compute this unrealized profit, we begin with the pool’s value (in USD), denoted as $x$, in the base token, and $owner\_share$, representing the pool owner’s portion of the pool’s asset. During the pool deployment phase, $x$ will be the owner’s initial deposit value: $x=x_{0}=Deposit_{0}$ and $owner\_share$ is 1, as they initially own 100% of the pool. For each subsequent $t^{th}$ DEX order made to the liquidity pool, the value of $x$ is updated as follows:
$x=x_{t}=x_{0}+\sum_{p=1}^{t}y_{p}$ (6)
@@
where $y_{p}$ represents the value of the base token amount (USD) for the $p^{th}$ DEX order. If the order is a buy or deposit (indicating a positive $y$), it signifies that $y$ USD worth of base token has been added to the pool through the order. Conversely, sell and withdraw orders indicate a decrease in value (indicating a negative $y$) due to the outflow of base tokens. Furthermore, for each “deposit” and “withdraw” order, the pool proportions shared amongst the liquidity providers change by,
$owner\_share_{t}=\begin{cases}\begin{array}[]{l}owner\_share_{t-1}\times\frac{x_{t-1}}{x_{t}}+\frac{y_{t}}{x_{t}}\\ \text{(if $t^{\text{th}}$ is a deposit/withdraw activity}\\ \text{from the owner),}\end{array}\\ \begin{array}[]{l}owner\_share_{t-1}\times\frac{x_{t-1}}{x_{t}}\\ \text{(if $t^{\text{th}}$ is a deposit/withdraw activity}\\ \text{from other investors).}\\ \end{array}\end{cases}$ (7)
@@
Eq. 7 indicate that ${owner\_share}_{t}$ increases if the owner deposits more to the pool, and conversely, it decreases if the owner executes withdraw orders. Similarly, the owner’s share will decrease if a user deposits into the pool and will increase if a user withdraws from it. The mechanism behind the owner pool’s share in Eq.7 is explained in more detail in D. Therefore, the liquidity pool’s owner’s first-month unrealized return is defined as follows:
@@
$\begin{split}\textit{Profit\_unrealized}_{1m}=\textit{Return\_unrealized}_{1m}\\ ={x}_{t}\times{owner\_share}_{t},\\ \text{where $\textit{Profit\_realized}_{1m}\geq 0$}\end{split}$ (8)
@@
where $t$ in this equation represents the order number executed in the pool on the DEX since its deployment to the end of the first month. If the liquidity pool has a positive realized profit threshold, indicating that it has reached the break-even point and already generated profit, we can consider the unrealized return as unrealized profit since it represents a profit beyond the realized gains.
Appendix D Owner’s Pool Share
@@
We explain how the owner’s liquidity share is updated or changed after a deposit or withdrawal in a DEX liquidity pool. We denote the share of owners at the $t^{th}$ order as $\mathcal{S}_{t}$. We present two scenarios to illustrate this process.
@@
    Deposit/withdraw excluding the onwer.  When a deposit or withdrawal order is executed in the pool and the sender is not the owner, the owner’s share is updated as follows:
    $\mathcal{S}_{t}=\mathcal{S}_{t-1}\times\frac{x_{t-1}}{x_{t}}.$
    The owner’s liquidity pool share before the $t^{th}$ order is calculated by multiplying their previous share $\mathcal{S}_{t-1}$ by the pool value before the order $x_{t-1}$. After the $t^{th}$ order, the updated share is determined by dividing the owner’s liquidity amount by the new pool value $x_{t}$.
@@
    Deposit/withdraw by the owner. When the pool owner executes the $t^{th}$ order with size $y_{t}$, we compute the new owner share as follows:
    $\mathcal{S}_{t}=\mathcal{S}_{t-1}\times\frac{x_{t-1}}{x_{t}}+\frac{y_{t}}{x_{t}}.$
    Whether the $t^{th}$ order involves a deposit or withdrawal, the ratio $\frac{y_{t}}{x_{t}}$ represents the percentage of the new liquidity part contributed by the owner relative to the pool’s value. This ratio, reflecting either an addition or subtraction, merges with the owner’s share prior to the $t^{th}$ order and updates their stake accordingly. If the $t^{th}$ order is a deposit by the owner, the ratio $\frac{y_{t}}{x_{t}}$ yields a positive value, as the pool value $x_{t}$ is inherently positive, thereby increasing the owner’s share post-update. Conversely, in the case of withdrawals, this ratio becomes negative, diminishing the owner’s share.
Appendix E Rationale for Seletected MLs
@@
    efficiency for real-time detection: the models are computationally efficient once trained. They can process new transactions quickly, ideal for (near-)real-time scenarios. They also scale well to high-throughput environments, such as mempool monitoring or transactional analysis, where rapid predictions on incoming data are critical.
@@
    compatibility with feature-based data: They excel in structured, tabular data environments, leveraging numerical and categorical features without requiring complex data transformations, unlike graph-based approaches (require transforming data into relational structures).
@@
    balance of complexity and accuracy: Random Forest and XGBoost capture non-linear patterns and interactions, while Logistic Regression provides a lightweight yet robust alternative for simpler cases, ensuring high performance with manageable complexity.
@@
Appendix F Additional Statistics
Table 2 presents the top 5 LPs with the highest profits. Table 3 shows the scammers with the most SLID LPs and total profits.
Table 2: LPs with the highest total profit generated
Pool address Paired Token Profit (USD)
@@ Appendix F Additional Statistics
Table 3: SLID scammers that deployed most SLID LPs
@@
Appendix G Data Collection and Features
Collected data. We report the statistics of collected data (Table 1) used in the data collection process (cf. §7.1).
Gathered features.  We show features (Table 4’s upper-side) from collected LPs and DEX order events from PoolData. We gather token security features for honeypot determination.
Extracted features in ML detectors.  We show 57 features (Table 4’s bottom-side) extracted from DEXData within a given time window $d$ for the training and testing activities of the ML detector. The features are categorized as:
@@ Appendix G Data Collection and Features
    Liquidity pool features (LPF). These illustrate the characteristics of LPs during the specified period, encompassing aspects of the pool’s lifetime, volume, and asset value.
@@
    Profit features (PF). Features in this category indicate the profit patterns generated by the liquidity pool for its owner, covering a wide range of metrics, including return, profit, and return on investment.
Appendix H Pseudocode
Our main program consists of three functions (Algorithm 2), each focusing on an aspect of the liquidity pools’ operations.

### v3

@@ VIII Conclusion
This paper introduces slow liquidity drain (SLID) scams as a stealthy and financially devastating threat to the DeFi ecosystem. By combining heuristic-based detection with machine learning, we empirically identified 3,117 SLID pools, exposing \$103.2M in cumulative losses. Our analysis focuses on Ethereum-based DEXs and the six largest platforms, which together account for 82.35% of total DEX volume. While this excludes smaller DEXs and other blockchains with different trading dynamics, the core SLID behaviors we identify are representative of dominant market activity and provide a strong foundation for future cross-chain analysis. We hope this study raises community awareness to prevent such daily losses. “Slow” does not mean safe; it silently drains your wallet.
Ethics Considerations
Privacy considerations. We focused our research on the behavior pattern of the owner that creates SLID liquidity pool and their activities. All examined data are publicly available on the Ethereum network, with all addresses anonymized by only keeping the first/last 5 characters of each, and generated mapping values for similarity determination (of determining transactions involving the same tokens/pools/users) in analysis tasks. We have no intention of deanonymizing any of the involved addresses for privacy reasons.
Respect for people. We prioritized user privacy throughout this research. All data analyzed was sourced from publicly available blockchain records, and no personally identifiable information (PII) was collected or processed. While blockchain data is inherently pseudonymous (also claimed before), we ensured that our findings respected user anonymity. Additionally, any insights derived from community-reported rug pulls were aggregated to avoid targeting individual users.
Appendix A Appendix
A-A Additional Related Work
@@
Liquidity pool, DEX and DeFi. Over the past five years, the rise of DeFi and DEXs has attracted millions of users and billions in inflows, leading to numerous research studies [80, 81, 82, 67, 9, 83, 84, 85]. Hendrik et al. [80] examined the opportunities and challenges for DeFi and its products in this mass adoption era. Fan et al. [82] conducted a comprehensive study on cryptocurrency trading activities for measuring market valuation. Jianlei et al. [67] investigated the DEXs roles and their impact on the development of the DeFi ecosystem. Saleh et al. [83] analyzed the effects of liquidity on Uniswap for liquidity providers and conducted estimations of fees and impermanent loss. Similarly, Lioba et al. [84] developed a theoretical model to analyze risks and returns faced by liquidity providers under various circumstances.
@@
Cryptocurrency scam detection. Various research efforts have focused on analyzing DEX scams, such as rug pull [24, 39, 86, 87] and honeypot [25, 56, 88]. For instance, Bruno et al. [24] utilized ML algorithms to identify rug pull and scam-likely tokens. Cernera et al. [39] conducted a longitudinal analysis of the lifetime of rug pull LPs on Ethereum and Binance Smart Chain. Rundong et al. [56] introduced various types of honeypot traps on DEXs along with their respective detection methods. Similarly, Christof et al. [25] conducted a comprehensive honeypot analysis with a proposed symbolic-execution-based solution. In addition, the scams originating from traditional finance, such as pump-and-dump schemes [89, 90] and Ponzi schemes [48, 91], have also been studied to provide warnings to investors. Several studies [92, 93, 94, 95] have suggested general fraud detection models that identify anomalous transactions or users through graph neural networks. However, among all the studies in this field, the SLID scam has yet to be discovered.
A-B LP Owner Financial Guarantee (cf. \$II)
@@
Let the LP’s initial deposit (by the owner) be $x$ units of the paired token and $y$ units of the base token. From the formula of the AMM model from Uniswap [35], the relationship between these tokens in the LP is reflected by:
@@
where $k$ is a constant determined by the initial liquidity and remains unchanged. With the mentioned setup, the pool’s owner initially monopolizes the paired token supply, which means no paired tokens $x$ exist outside the LP before the pool’s deployment. Any investor/user that is investing/interacting with the LP must first buy the paired token from the LP by swapping it with base tokens. Let $\Delta x$ represent the total amount of the paired token bought by users or distributed as incentives (for users that deposit a pair of tokens for liquidity providers). The formula will be rewritten to represent the pool’s state at any moment after deployment:
@@
where $\Delta y$ is the change of base token amount in the LP due to the change of paired token, and x’/y’ is the amount of paired/base token in the LP after the change, compared to the initial deposit amount (from the owner). Under the AMM model, the constant $k$ must not change.
In the worst scenario where no investor interacts with the LP ($\Delta x=0$) or all investors sell back their paired token ($\Delta x$, since the monopoly of paired tokens from the owner made the limit of paired token that investor possess totally at any time cannot be larger than $\Delta x$) to the pool, the base token possessed by the owner afterward is computed by:
$y^{\prime}=\frac{k}{x^{\prime}}=\frac{x\times y}{x+\Delta x-\Delta x}=\frac{x}{x}\times y=y$ (4)
This proves that in the worst-case scenario, where the LP has no investor or all investors that previously bought (and used the bought paired token to deposit for incentives) sell back all their received paired tokens to the pool, the base token balance controlled by the owner will still be equal to $y$. Hence, with this setup, the LP owner is guaranteed to maintain their original investment of base tokens $y$, which contain real value, under any sequence of investor interactions after deployment.
A-C Computing Equation
@@
$\begin{split}\textit{Profit\_realized}=&\left(\sum_{i=1}^{m}Sell^{i}+\sum_{j=1}^{n}Withdraw^{j}\right)\\ &-\left(\sum_{k=1}^{o}Buy^{k}+\sum_{l=1}^{p}Deposit^{l}+\text{GAS}\right)\end{split}$ (5)
@@
where the owner’s investment activity (which adds assets in base tokes to the pool) includes $Buy^{k}$ and $Deposit^{l}$, representing the asset value (in USD) added to their LP during the $k^{\text{th}}$ buy and $l^{\text{th}}$ deposit orders, respectively. In contrast, the owner’s realized return consists of $Sell^{i}$ and $Withdraw^{j}$, which are the asset values in base token (in USD) that the owner removed from the LP during the $i^{\text{th}}$ sell and $j^{\text{th}}$ withdraw orders, respectively. The metric also accounts for the total network gas fee (GAS) used for LP deployment and all DEX activities performed by the owner. Since the equation focused on computing realized profit, all of the variables are only around the owner’s balance change of the base token (which holds the real value) through their activity in the LP (which also involves paired tokens).
@@
Unrealized profit computing equation. To compute this unrealized profit, we begin with the pool’s value (in USD), denoted as $x$, in the base token, and owner_shareowner\_share, representing the LP owner’s portion of the pool’s asset. During the LP deployment phase, $x$ will be the owner’s initial deposit value: $x=x_{0}=Deposit_{0}$ and owner_shareowner\_share is 1, as they initially own 100% of the pool. For each subsequent $t^{th}$ DEX order made to the LP, the value of $x$ is updated as follows:
$x=x^{t}=x_{0}+\sum_{p=1}^{t}y^{p}$ (6)
@@
where $y^{p}$ represents the value of base tokens (USD) for the $p^{th}$ DEX order. If the order is a buy or deposit (indicating a positive $y$), it signifies that $y$ USD worth of base token has been added to the LP through the order. Conversely, sell and withdraw orders indicate a decrease in value (indicating a negative $y$) due to the outflow of base tokens. Further, for each “deposit” and “withdraw” order, the LP proportions shared amongst the liquidity providers change by,
[…]
@@
Eq. 7 indicates that ${owner\_share}^{t}$ increases if the owner deposits more into the pool, and conversely, it decreases if the owner executes withdraw orders. Similarly, the owner’s share will decrease if a user deposits into the LP and will increase if a user withdraws from it. The mechanism behind the owner pool’s share in Eq. 7 is explained in more detail in §A-D. Therefore, the LP’s owner’s first-month unrealized return is defined as follows:
@@
$\begin{split}\textit{Profit\_unrealized}_{1m}=\textit{Return\_unrealized}_{1m}\\ ={x}^{t}\times{owner\_share}^{t},\\ \text{where $\textit{Profit\_realized}_{1m}\geq 0$}\end{split}$ (8)
@@
where $t$ in this equation represents the order number executed in the LP on the DEX since its deployment to the end of the first month. If the LP has a positive realized profit threshold (having reached the break-even point and already generated profits), we can consider the unrealized return as unrealized profit since it represents a profit beyond the realized gains.
A-D Owner’s LP Share
@@
We explain how the owner’s liquidity share is updated after a deposit or withdrawal in a DEX LP. We denote the share of owners at the $t^{th}$ order as $owner\_share^{t}$. We present two scenarios.
@@
    Deposit/withdraw excluding the owner. When a deposit or withdrawal order is executed in the LP and the sender is not the owner, the owner’s share is updated as follows:
    $owner\_share^{t}=owner\_share^{t-1}\times\frac{x^{t-1}}{x^{t}}.$
    The owner’s LP share before the $t^{th}$ order is calculated by multiplying their previous share $owner\_share^{t-1}$ by the LP value before the order $x^{t-1}$. After the $t^{th}$ order, the updated share is determined by dividing the owner’s liquidity amount by the new LP value $x^{t}$.
@@
    Deposit/withdraw by the owner. When the LP owner executes the $t^{th}$ order with size $y^{t}$, we compute the new owner share as follows:
    $owner\_share^{t}=owner\_share^{t-1}\times\frac{x^{t-1}}{x^{t}}+\frac{y^{t}}{x^{t}}.$
    Whether the $t^{th}$ order involves a deposit or withdrawal, the ratio $\frac{y^{t}}{x^{t}}$ represents the percentage of the new liquidity part contributed by the owner relative to the pool’s value. This ratio, reflecting either an addition or subtraction, merges with the owner’s share prior to the $t^{th}$ order and updates their stake accordingly. If the $t^{th}$ order is a deposit by the owner, the ratio $\frac{y^{t}}{x^{t}}$ yields a positive value, as the LP value $x^{t}$ is inherently positive, thereby increasing the owner’s share post-update. Conversely, in the case of withdrawals, this ratio becomes negative, diminishing the owner’s share.
A-E Dataset Reliability Evaluation (cf. \$VI)
We also evaluated the reliability of our collected ground truth that was labeled from the data collection process.
  • (manual observation) We selected 100 SLID pools randomly and our analysis revealed no sign of false positives. For every pool we reviewed, we first verified it with well‑established scam traits: deployers ($i$) still held the full proportion of LP tokens; ($ii$) executed a sequence of profit-taking sequence that stepped the price down; and (iiiiii) had never burned liquidity. Next, we cross‑checked these pools against our declarative filters to guarantee all heuristic signals are detected in each flagged LPs. Our manual review found zero false positives where all 100 pools satisfied scam traits and heuristic conditions.
  • (external cross-check) We tested the 100 randomly selected SLID pools using existing honeypot and rug pull heuristics, none of which detected the scam. Moreover, we randomly collected 20 verified rug‑pull tokens from ChainAbuse [41] and 20 verified honeypots from GoPlus [47] and cross-matched their contract addresses with our 3,117 SLID labels. The intersection was empty (both false positive/negative =0=0), indicating perfect precision and recall (=1=1) and showing that our filter does not misclassify other scam types as SLID.
A-F The Role of ML in SLID Detection (cf. §VI-C)
A-F1 Why ML is essential for our SLID detection
We have two main reasons as below.
(ii) Early detection when SLID pools fail to form required characteristics. In the rule-based heuristic, certain features indicative of SLID scams take time to form, as the scam must exhibit multiple defining characteristics before it can be detected through our fuzzy heuristic. These features, such as gradual liquidity withdrawal, sales of inflated tokens, or lifetime, are often subtle at first and accumulate over time. As a result, heuristic methods may struggle to identify SLID scams in their early stages.
Machine learning offers an advantage by detecting early-stage SLID cases that have not yet fully formed with required characteristics. ML models are capable of recognizing patterns in data that may seem insignificant to human-defined rules, such as early signs of unusual behavior, even when those behaviors are not yet obvious. The early detection capability was validated in our experiments, where the ML model’s prediction time was 4.77 times faster than the heuristic approach (Figure 13).
(iiii) Detecting borderline cases. ML can be used for detecting borderline cases or characteristics by adjusting its threshold. In this study, we define a borderline SLID pool as an LP that satisfies only a subset of the formal SLID conditions (Definition 1, §V-A), such as:
  • $\bullet$
    The pool’s owner is taking profit but less than 5 times ($1\leq T_{count}<5$).
  • $\bullet$
    The pool yielded zero profit in the first month, but became positive and continues to grow in subsequent periods.
Since this paper sets out the first evidence‑based foundation for SLID detection, every filter in the heuristic of Definition 1 (§V-A) is served with a fixed range of thresholds (based on the analysis from §IV) to minimize false positives. As a result, such borderline pools continue to pose risks to liquidity providers, given that their subtle or noisier behavioral patterns tend to evade detection by rigid rule-based methods.
Unlike rule-based heuristics, which rely on rigid thresholds and predefined patterns (although we’ve relaxed the thresholds for more real-time capturing enhancement), ML approaches can implicitly incorporate uncertainty and behavioral noise [96, 97]. They excel at identifying patterns in data that may be too complex or nuanced. By leveraging training on large, diverse datasets, ML models can detect these early, emerging signs of SLID scams even before the fraudulent activity becomes fully apparent or fits within a fixed heuristic definition. Multiple previous studies on DeFi scams have successfully employed machine learning techniques to detect borderline cases [49, 91, 24].
A-F2 Rationale for selected MLs
@@
    efficiency for real-time detection: Those models process transactions quickly and are computationally efficient once trained, ideal for (near-)real-time scenarios. They also scale well to high-throughput cases (e.g., memLP monitoring, transactional analysis), where rapid predictions on incoming data are critical.
@@
    compatibility with feature-based data: They excel in structured, tabular data environments, leveraging numerical and categorical features without requiring complex data transformations, unlike graph-based approaches (which require transforming data into relational structures).
@@
    balance of complexity and accuracy: Random Forest and XGBoost capture non-linear patterns. Logistic Regression provides a lightweight yet robust alternative for simpler cases, with high performance and manageable complexity.
@@
A-G Additional Statistics (cf. §VI-B)
Table II presents the top 5 LPs with the highest profits. Table III shows the scammers with the most SLID LPs.
Table II: LPs with the highest total profit generated
LP address Paired Token Profit (USD)
@@
Table III: SLID scammers that deployed most SLID LPs
@@
A-H Data Collection and Features (cf. §VI-A)
Collected data. We report the statistics of collected data (Table I) used in the data collection process (cf. §VI-A).
Gathered features. We show features (Table IV’s upper-side) from collected LPs and DEX order events from PoolData. We gather token security features for honeypot determination.
Extracted features in ML detectors. We show 57 features (Table IV’s bottom-side) extracted from DEXData within a given time window $d$ for the training and testing activities of the ML detector. The features are categorized as:
@@
    LP features (LPF). These illustrate the characteristics of LPs during the specified period, encompassing aspects of the pool’s lifetime, volume, and asset value.
@@
    Profit features (PF). They indicate the profit patterns generated by LP for its owner, covering a wide range of metrics, including return, profit, and return on investment.
A-I Pseudocode
Our main program consists of three functions (Algorithm 2), each focusing on an aspect of the LPs’ operations.